An alpha particle moves toward a rest nucleus if k.e is 10 MeV and atomic no of nucleus is 50.the closest approach will be

Grade:12

2 Answers

Nandakumar
6 years ago
I think you have to equate kinetic and potential energy ie 1/2mv2=1/4pi€×2Ze2/r r is the distance of closest approach by substituting the values you get r=14.4×10-15 m
Pratham Savjani
4 years ago
Here we need to equate Potential Energy and Kinetic energy.
Therefore
K.E=P.E
1/2 mv^2 = kq1q2/r
r=kq1q2/0.5mv^2
r=9*10^9*2e*50e/10 mev
Search the value of mev in terms of joules and substitute it.
e=1.6*10^-19
You will get answer as 1.44*10^-14